OpenStudy (anonymous):

Suppose we have the radical function, f(x)=(7-∜(3x+1))/5 What is the domain of f? Compute f(5)-5

OpenStudy (anonymous):

f(5) = [7 - (3(5) + 1) ^(1/4) ] -5 f(5) = [7 - 2]/ 5 f(5) = 5/5 = 1 Thus, f(5) - 5 = -4

OpenStudy (anonymous):

then what would the domain be?

OpenStudy (anonymous):

all real numbers?

OpenStudy (anonymous):

domain you have so solve \[3x+1\geq 0\] \[3x\geq -1\] \[ x\geq -\frac{1}{3}\]

OpenStudy (anonymous):

I thought that the domain was suppose to be the numbers that don't make the denominator 0...

OpenStudy (anonymous):

when you take an even root of a negative number, you get a nonreal result. Therefore; when you take an even root, any x that makes the radicand negative is excluded from the domain

OpenStudy (anonymous):

ah...okay. Thanks guys :)